OpenStudy (anonymous):

Solve: 3c + 2 = 4c - 9 Solve: 3c + 2 = 4c - 9 @Mathematics

OpenStudy (aroub):

-4c+3c=-2-9 -c=-11 c=11

OpenStudy (anonymous):

If you don't mind me asking, where'd you get the negative 4 from? :)

OpenStudy (anonymous):

it's from moving the 4c in '4c-9' across the sign.. 3c+2=4c-9 -4c -4c This is to get all variables on the same side of the equals sign

OpenStudy (anonymous):

to work it out... 3c+2=4c-9 is your original problem. You need to get variables and numbers each on their own side, so.. 3c+2=4c-9 -4c -4c to get -1c+2=-9. Then you would... -2 -2 to get -1c = -11 then divide both sides by -1 to get c = 11